# Non-Stationary Covariance Functions for Spatial Data on Linear Networks

### v1

@@ 3.2 Examples
A direct calculation shows that $K(s,t)$ in Equation (3) can be written as
@@
$\psi(r)=\int_{0}^{\infty}\frac{1}{\sqrt{r+w}}\text{d}F(w),\qquad r>0.$ (8)
Functions of the form (8) belong to the class of generalized Stieltjes functions of order $1/2$ (see, e.g., Koumandos and Pedersen, 2019). When $a$ and $b$ are constant functions, we recover a subclass of the isotropic covariance functions presented in Anderes et al. (2020), since every generalized Stieltjes function is completely monotone (see, e.g., Menegatto, 2020, Section 2.2).
Since closed-form expressions for generalized Stieltjes functions are available in standard tables of integral transforms such as (Erdélyi, 1954, Chapter 14), it is straightforward to derive explicit examples of non-stationary covariance functions.
@@
Let $F$ be the Dirac delta measure at $\beta>0$, that is, $\text{d}F(w)=\text{d}\delta_{\beta}(w)$. Then, we immediately obtain the covariance function
$K(s,t)=\sqrt{\frac{b(s)b(t)}{\alpha(s,t)}}\left(1+\frac{\beta\text{d}_{R}(s,t)}{\alpha(s,t)}\right)^{-1/2}.$ (9)
@@
Let $\text{d}F(w)=(1/\beta)\exp(-w/\beta)\text{d}w$, for $\beta>0$. Then, using Equation 14.4.10 in Erdélyi (1954), we obtain
$K(s,t)=\begin{cases}\displaystyle\frac{b(s)}{\sqrt{a(s)}},&\text{ if }s=t,\\ \\ \displaystyle\sqrt{\frac{\pi b(s)b(t)}{\beta}}\exp\left(\frac{\alpha(s,t)}{\beta\text{d}_{R}(s,t)}\right)\text{erfc}\left(\sqrt{\frac{\alpha(s,t)}{\beta\text{d}_{R}(s,t)}}\right),&\text{ if }s\neq t,\end{cases}$
@@
Let $\text{d}F(w)=\sqrt{\beta}\pi^{-1}w^{-1/2}(\beta+w)^{-1}\text{d}w$, for $\beta>0$. This is a scaled Beta prime probability measure. Then, using Equation 14.4.9 in Erdélyi (1954), we obtain
$K(s,t)=\begin{cases}\displaystyle\frac{b(s)}{\sqrt{a(s)}},&\text{ if }s=t,\\ \\ \displaystyle\frac{2}{\pi}\sqrt{\frac{b(s)b(t)}{\beta}}\,_{2}F_{1}\left(1,1/2,3/2;1-\frac{\alpha(s,t)}{\beta\text{d}_{R}(s,t)}\right),&\text{ if }s\neq t,\end{cases}$
@@ 3.3 Illustration
To illustrate our proposal, we simulate a GRF on a linear network formed by the mortar lines of a brick wall. This network, consisting of 156 vertices and 203 edges, and available in the R package spatstat (Baddeley and Turner, 2005), has been the focus of studies on the positions of spider webs of the urban wall spider Oecobius navus (Voss et al., 2007).
We use the covariance function in Equation (9) with $\beta=1$ and the parameterization $b(s)=c(s)\sqrt{a(s)}$, for some function $c:\mathcal{L}\rightarrow(0,\infty)$, which ensures that ${\rm var}[X(s)]=K(s,s)=c(s)$, for all $s\in\mathcal{L}$. Specifically, we consider
$a(s)=2500\times\exp\left(-\frac{\text{d}_{R}(s,s^{*})}{100}\right),\qquad s\in\mathcal{L},$ (10)
@@
$c(s)=2\times\exp\left(-\frac{\text{d}_{R}(s,s^{*})}{1000}\right),\qquad s\in\mathcal{L},$ (11)
@@
where $s^{*}$ is a fixed point on $\mathcal{L}$. Both functions attain their maximum at $s^{*}$ and decay exponentially as $s$ moves away from this point, as illustrated in Figure 1. Accordingly, we expect the simulated GRF to exhibit larger correlation range and higher variance at locations near $s^{*}$, decreasing as one moves away from this point.
@@
Figure 1: Functions $a$ (left) and $c$ (right), defined in Equations (10) and (11), respectively, on the spiders linear network. The circle indicates the point $s^{*}$ involved in the definitions of $a$ and $c$.
@@
To better illustrate the local behavior of this covariance function, Figure 2 shows the curves $\text{d}_{R}\mapsto\sqrt{a_{0}}/\sqrt{a_{0}+\text{d}_{R}}$ for different values of $a_{0}$. These curves capture part of the local isotropic correlation structures that the covariance function exhibits. They are displayed over the interval $(0,800)$ since the maximum resistance distance between two points in this domain is approximately 837.
@@
Figure 2: Mapping $\text{d}_{R}\mapsto\sqrt{a_{0}}/\sqrt{a_{0}+\text{d}_{R}}$ for selected values of $a_{0}$, illustrating how the local isotropic correlation structure depends on $a_{0}$.
@@
Figure 3 (left panel) shows the simulated GRF over 5,684 sites (28 per edge). The field attains high values at sites in the top part of the linear network, and because the correlation range is greater in this region, these high values extend over a wide area. In contrast, at sites in the lower part of the linear network, the field varies more rapidly, reflecting a smaller correlation range.
@@
Figure 3: (Left) Simulated GRF on the spiders linear network, with the non-stationary covariance function given in Equation (9). (Right) Corresponding realization of a LGCP, with line thickness proportional to the magnitude of the log Gaussian field.
@@
The right panel of Figure 3 shows the realization of the corresponding LGCP, where the log Gaussian random field is defined as $\Lambda(s)=0.002\times\exp\{X(s)\}$, with $X(s)$ denoting the underlying GRF. Recall that, conditioned on $\Lambda$, this point process on $\mathcal{L}$ is almost surely a Poisson process with intensity function $\Lambda$ (with respect to the arc length measure on $\mathcal{L}$). As expected, more points occur in the upper part of the network, covering a larger area, while fewer points are present in the lower part.
For a more detailed treatment of Cox processes on linear networks, we refer the reader to Møller and Rasmussen (2024).
@@ 4.1 Description of the method
We consider the weighted local likelihood estimation proposed by Anderes and Stein (2011). We now describe the method and the necessary modifications required to adapt it to the network context.
Suppose that we observe a realization of a GRF $\{X(s):s\in\mathcal{L}\}$ at locations $s_{1}\ldots,s_{N}$ in $\mathcal{L}$. For any given site $s\in\mathcal{L}$, let $\mathcal{N}_{s,k}$ denote the set of $k$ observations nearest to $s$ with respect to the resistance metric. Recall that our goal is to estimate $a$ and $b$ at the selected site $s$. Given weights $\lambda_{1},\ldots,\lambda_{N}$, the weighted local likelihood function is defined as
@@
where $\text{Lik}(a_{0},b_{0}\mid\mathcal{N}_{s,k})$ denotes the log-likelihood of the data in $\mathcal{N}_{s,k}$, with $a_{0}$ and $b_{0}$ representing fixed values of the functions $a(\cdot)$ and $b(\cdot)$. As in Anderes and Stein (2011), we define $\text{Lik}(a_{0},b_{0}\mid\mathcal{N}_{s,0})$ as 0.
@@
The principle behind this method is that we locally fit an isotropic model, not simply by taking a window of data around $s$ (a strategy also used in Anderes and Stein, 2011, which has been shown to be less competitive), but by considering inflating neighborhoods and gradually assigning less weight to those that cover a very large region. For this reason, the weights include a bandwidth parameter to control this effect.
4.2 Simulation study
@@
$\lambda_{k}^{\text{cs}}=\left(1-\frac{\text{d}_{R}(s,s_{k})}{\tau}\right)_{+},\qquad k=1,\ldots,N,$ (12)
@@
$\lambda_{k}^{\text{g}}=\exp\left(-\frac{\text{d}_{R}(s,s_{k})^{2}}{\eta^{2}}\right),\qquad k=1,\ldots,N,$ (13)
@@
We consider the Chicago street network available in the R package spatstat, and the covariance function in Equation (9), with the parameterization discussed in Section 3.3. Following Anderes and Stein (2011), we assume for simplicity that the variance of the random field is known and equal to one, so that we focus on the estimation of the local practical range. The function $a$ used in our study is shown in the left panel of Figure 4. We also highlight six randomly selected locations on the network at which $a$ will be estimated.
We simulate 100 independent realizations of the GRF on this domain at 503 locations (one at the midpoint of each edge). For each realization, we estimate $a$ at the six selected sites. The weighted local likelihood method is implemented with the weighting schemes described above. We use bandwidth parameters $\eta=120$ and $\tau=200$, which produce weight functions with a comparable spatial extent but distinct structural features (see the right panel of Figure 4). For reference, the maximum resistance distance between two points in this network is approximately 675.
Figure 4: (Left) The function $a$ used in our simulations on the Chicago street network. Six randomly selected locations where $a$ will be estimated are highlighted. (Right) Decay of the weights as a function of the resistance metric for the compactly supported ($\tau=120$) and Gaussian ($\eta=120$) weighting schemes.
In Figure 5, we show the centered boxplots of the estimates across the 100 realizations, for each weighting scheme and each of the six selected sites. Overall, no significant bias is observed in any of the boxplots, demonstrating the effectiveness of the method. One point of interest is the higher variance in the estimates at site 2. This effect is consistent with the site’s proximity to the boundary of the network, which reduces the number of neighbors that effectively contributes to the local estimate. A similar observation was reported by Anderes and Stein (2011).
Figure 5: Centered boxplots of the estimates of $a$, at the six selected sites, across 100 independent realizations of a GRF with covariance function (9). Results are shown for the compactly supported and Gaussian weighting schemes.
@@
We also compare the computational performance of the weighting schemes by measuring evaluation times as a function of the sample size of the GRF. Specifically, we simulate GRFs with $N=503,1006,1509$, and 2012, which correspond to taking 1, 2, 3, and 4 locations per edge, respectively, and implement the method using compactly supported weights with bandwidth parameters $\tau=100,150$, and 200. For each configuration, we report the time required for a single evaluation of the weighted local likelihood function. The results are summarized in Table 1, allowing us to assess the computational gains relative to the Gaussian weighting scheme. For $\tau=100$ and 150, the reduction in computation time is particularly pronounced. For instance, when $\tau=150$, the computation time is reduced by approximately $50\%$ compared with Gaussian weighting scheme, whereas for $\tau=100$, evaluating the objective function takes at most $2\%$ of the time required with Gaussian weights, and for most sample sizes the percentage is even lower.
Table 1: Times (in seconds) for a single evaluation of the weighted local likelihood function under each weighting scheme, as a function of the sample size NN. For the compactly supported weights we consider $\tau=100,150$, and 200. The simulation was performed on an Apple M1 chip with 8GB of memory.
@@
Our findings both support the observations reported in Anderes and Stein (2011) regarding reduced precision near the domain boundaries and provide additional insights into the statistical and computational performance of different weighting schemes in our specific setting. A promising strategy is to use a combination of weights, employing compactly supported weights in the interior and adapting the weights near the boundaries. We refer the reader to Anderes and Stein (2011) for a more comprehensive treatment of weights and bandwidth selection.

### v2

@@ 3.2 Examples
A direct calculation shows that the covariance function in Equation (3.1) admits the following representation:
@@
$\psi(r)=\int_{0}^{\infty}\frac{1}{\sqrt{r+w}}\text{d}F(w),\qquad r>0.$ (3.2)
Functions of the form (3.2) belong to the class of generalized Stieltjes functions of order $1/2$ (see, e.g., Koumandos and Pedersen,, 2019). When $a$ and $b$ are constant functions, we recover a subclass of the isotropic covariance functions presented in Anderes et al., (2020), since every generalized Stieltjes function is completely monotone (see, e.g., Menegatto,, 2020, Section 2.2).
Since closed-form expressions for generalized Stieltjes functions are available in standard tables of integral transforms such as (Erdélyi,, 1954, Chapter 14), it is straightforward to derive explicit examples of non-stationary covariance functions.
@@
Let $F$ be the Dirac delta measure at $\eta>0$, that is, $\text{d}F(w)=\text{d}\delta_{\eta}(w)$. Then, we immediately obtain the covariance function
$K(s,t)=\sqrt{\frac{b(s)b(t)}{\alpha(s,t)}}\left(1+\frac{\eta\text{d}_{R}(s,t)}{\alpha(s,t)}\right)^{-1/2}.$ (3.3)
@@
Let $\text{d}F(w)=(1/\eta)\exp(-w/\eta)\text{d}w$, for $\eta>0$. Then, using Equation 14.4.10 in Erdélyi, (1954), we obtain
$K(s,t)=\begin{cases}\displaystyle\frac{b(s)}{\sqrt{a(s)}},&\text{ if }s=t,\\ \\ \displaystyle\sqrt{\frac{\pi b(s)b(t)}{\eta}}\exp\left(\frac{\alpha(s,t)}{\eta\text{d}_{R}(s,t)}\right)\text{erfc}\left(\sqrt{\frac{\alpha(s,t)}{\eta\text{d}_{R}(s,t)}}\right),&\text{ if }s\neq t,\end{cases}$
@@
Let $\text{d}F(w)=\sqrt{\eta}\pi^{-1}w^{-1/2}(\eta+w)^{-1}\text{d}w$, for $\eta>0$. This is a scaled Beta prime probability measure. Then, using Equation 14.4.9 in Erdélyi, (1954), we obtain
$K(s,t)=\begin{cases}\displaystyle\frac{b(s)}{\sqrt{a(s)}},&\text{ if }s=t,\\ \\ \displaystyle\frac{2}{\pi}\sqrt{\frac{b(s)b(t)}{\eta}}\,_{2}F_{1}\left(1,1/2,3/2;1-\frac{\alpha(s,t)}{\eta\text{d}_{R}(s,t)}\right),&\text{ if }s\neq t,\end{cases}$
@@ 3.3 Illustration
To illustrate our proposal, we simulate a GRF on a linear network formed by the mortar lines of a brick wall. This network, consisting of 156 vertices and 203 edges, and available in the R package spatstat (Baddeley and Turner,, 2005), has been the focus of studies on the positions of spider webs of the urban wall spider Oecobius navus (Voss et al.,, 2007).
We use the covariance function in Equation (3.3) with $\eta=1$ and the parameterization $b(s)=c(s)\sqrt{a(s)}$, for some function $c:\mathcal{L}\rightarrow(0,\infty)$, which ensures that ${\rm var}[X(s)]=K(s,s)=c(s)$, for all $s\in\mathcal{L}$. Specifically, we consider
$a(s)=2500\times\exp\left(-\frac{\text{d}_{R}(s,s^{*})}{100}\right),\qquad s\in\mathcal{L},$ (3.4)
@@
$c(s)=2\times\exp\left(-\frac{\text{d}_{R}(s,s^{*})}{1000}\right),\qquad s\in\mathcal{L},$ (3.5)
@@
where $s^{*}$ is a fixed point on $\mathcal{L}$. Both functions attain their maximum at $s^{*}$ and decay exponentially as $s$ moves away from this point, as illustrated in Figure 3.1. Accordingly, we expect the simulated GRF to exhibit larger correlation range and higher variance at locations near $s^{*}$, decreasing as one moves away from this point.
@@
Figure 3.1: Functions $a$ (left) and $c$ (right), defined in Equations (3.4) and (3.5), respectively, on the spiders linear network. The circle indicates the point $s^{*}\in\mathcal{L}$ involved in the definitions of $a$ and $c$.
To better illustrate the local behavior of this covariance function, Figure 3.2 shows the curves
$\text{d}_{R}\mapsto\frac{\sqrt{a_{0}}}{\sqrt{a_{0}+\text{d}_{R}}}$
@@
for different values of $a_{0}$. These curves capture part of the local isotropic correlation structures that the non-stationary covariance function exhibits. They are displayed over the interval $(0,800)$ since the maximum resistance distance between two points in this domain is approximately 837.
@@
Figure 3.2: Mapping $\text{d}_{R}\mapsto\sqrt{a_{0}}/\sqrt{a_{0}+\text{d}_{R}}$ for selected values of $a_{0}$, illustrating how the local isotropic correlation structure of the non-stationary covariance function (3.3) depends on $a_{0}$.
@@
Figure 3.3 (left panel) shows the simulated GRF over 5,684 sites (28 per edge). The simulated field attains high values at sites in the top part of the linear network, and because the correlation range is greater in this region, these high values extend over a wide area. In contrast, at sites in the lower part of the linear network, the simulated field varies more rapidly, reflecting a smaller correlation range.
@@
Figure 3.3: (Left) Simulated GRF on the spiders linear network, with the non-stationary covariance function given in Equation (3.3). (Right) Corresponding realization of a LGCP, with line thickness proportional to the magnitude of the log Gaussian field.
@@
The right panel of Figure 3.3 shows the realization of the corresponding LGCP, where the log Gaussian random field is defined as $\Lambda(s)=0.002\times\exp\{X(s)\}$, with $X(s)$ denoting the underlying GRF. Recall that, conditioned on $\Lambda$, this point process on $\mathcal{L}$ is almost surely a Poisson process with intensity function $\Lambda$ (with respect to the arc length measure on $\mathcal{L}$). As expected, more points occur in the upper part of the network, covering a larger area, while fewer points are present in the lower part. For a more detailed treatment of Cox processes on linear networks, we refer the reader to Møller and Rasmussen, (2024).
4 Extensions of the framework
4.1 Stochastic representation
The following proposition provides an explicit construction of an elementary random field whose covariance function is given by (3.1). Throughout this subsection, we assume without loss of generality that $F$ is a probability measure on $(0,\infty)$.
Proposition 4.1.
Let $Z$ be a real-valued random variable with probability density function $g$. Let $W$ be a positive random variable with distribution $F$. Let $\epsilon$ be a random variable with zero mean and unit variance. Let $\{A(s):s\in\mathcal{L}\}$ denote the auxiliary GRF defined in Equation (2.1). Assume that $Z$, $W$, $\epsilon$, and $A$ are mutually independent. Let $\phi(\cdot;\mu,\sigma^{2})$ denote the probability density function of a normal distribution with mean $\mu$ and variance $\sigma^{2}$. Then, the random field
$X(s)=\epsilon\,\sqrt{\frac{b(s)}{g(Z)}}\left(\frac{\pi}{W}\right)^{1/4}\phi\left(Z;\frac{A(s)}{\sqrt{2}},\frac{a(s)}{4W}\right),\qquad s\in\mathcal{L},$ (4.1)
has zero mean and covariance function given by (3.1).
A proof of this result is provided in Appendix B.
The stochastic representation introduced in Equation (4.1) extends Proposition 3.4 of Alegría et al., (2026) from the isotropic setting to the non-stationary case.
The importance of the stochastic representation (4.1) lies in its potential for developing computationally efficient algorithms to simulate GRFs with covariance functions of the form (3.1). Recall that simulation based on the Cholesky decomposition of the covariance matrix has cubic computational complexity, which motivates the development of more efficient methods (see Lantuéjoul,, 2001 for a comprehensive treatment in Euclidean spaces).
Since the random field in (4.1) is not Gaussian, one strategy is to add several independent copies of this elementary random field, rescale appropriately, and then invoke a central limit approximation. More precisely, an approximately GRF with covariance function (3.1) is given by
$X(s)=\frac{1}{\sqrt{J}}\sum_{j=1}^{J}\epsilon_{j}\,\sqrt{\frac{b(s)}{g(Z_{j})}}\left(\frac{\pi}{W_{j}}\right)^{1/4}\phi\left(Z_{j};\frac{A^{(j)}(s)}{\sqrt{2}},\frac{a(s)}{4W_{j}}\right),\qquad s\in\mathcal{L},$
where $J$ is a large positive integer, and $\epsilon_{1},\ldots,\epsilon_{J}$, $Z_{1},\ldots,Z_{J}$, $W_{1},\ldots,W_{J}$, and $A^{(1)},\ldots,A^{(J)}$ are independent copies of the random ingredients involved in Proposition 4.1.
We refer the reader to (Alegría et al.,, 2026, Section 4.5) for a more detailed treatment of the computational aspects of these types of algorithms in the linear network scenario, as well as for implementation guidelines.
4.2 Vector-valued GRFs
In multivariate spatial statistics, several georeferenced variables are analyzed simultaneously (Wackernagel,, 2003). Interest focuses not only on the marginal covariance functions of each variable but also on the cross-covariance functions between them.
Formally, consider a vector-valued GRF, namely $\{\bm{X}(s)=(X_{1}(s),\ldots,X_{q}(s))^{\top}:s\in\mathcal{L}\}$, where $q$ is the number of components, and ${\rm var}[X_{i}(s)]<\infty$, for all $s\in\mathcal{L}$ and $i=1,\ldots,q$. In this setting, the covariance function is matrix-valued, $\bm{K}:\mathcal{L}\times\mathcal{L}\rightarrow\mathbb{R}^{q\times q}$, with entries
$K_{ij}(s,t)={\rm cov}\big[X_{i}(s),X_{j}(t)\big],\qquad s,t\in\mathcal{L},\,\,i,j=1,\ldots,q.$
The covariance function must be positive semidefinite in the sense that, for each $N\in\mathbb{N}$ and any collection of locations $s_{1},\ldots,s_{N}\in\mathcal{L}$, the covariance matrix of the $Nq$-dimensional vector $(\bm{X}(s_{1})^{\top},\ldots,\bm{X}(s_{N})^{\top})^{\top}$ is positive semidefinite.
The following theorem provides a construction of non-stationary matrix-valued covariance functions, extending Theorem 3.1 to the vector-valued setting.
Theorem 4.1.
Let $a_{i},b_{i}:\mathcal{L}\rightarrow(0,\infty)$, for $i=1,\ldots,q$, and define $\alpha_{ij}(s,t)=[a_{i}(s)+a_{j}(t)]/2$. Let $\bm{\beta}=[\beta_{ij}]_{i,j=1}^{q}$ be a symmetric positive semidefinite matrix. Let $\text{d}_{R}(s,t)$ denote the resistance metric between $s$ and $t$, and let $F$ be a positive finite measure on $(0,\infty)$. Then, the matrix-valued function $\bm{K}:\mathcal{L}\times\mathcal{L}\rightarrow\mathbb{R}^{q\times q}$, with entries
$K_{ij}(s,t)=\displaystyle\int_{0}^{\infty}\frac{\beta_{ij}\sqrt{b_{i}(s)b_{j}(t)}}{\sqrt{\alpha_{ij}(s,t)+\text{d}_{R}(s,t)w}}\text{d}F(w),\qquad s,t\in\mathcal{L},\,\,i,j=1,\ldots,q,$ (4.2)
is positive semidefinite.
In Appendix C, we provide a proof of Theorem 4.1. This theorem is the network counterpart of a result presented in Kleiber and Nychka, (2012) for Euclidean spaces.
In this construction, the functions $b_{i}$ and $a_{i}$ allow us to control, respectively, the local variance and correlation range of the $i$-th component of the vector-valued GRF, in a similar fashion to the univariate ($q=1$) case. The parameters $\beta_{ij}$, in turn, allow us to modulate the colocated correlation coefficient between the $i$-th and $j$-th components. More precisely, the effect of $\beta_{ij}$ on the colocated correlation is given by
${\rm cor}[X_{i}(s),X_{j}(s)]=\frac{{\rm cov}[X_{i}(s),X_{j}(s)]}{\sqrt{{\rm var}[X_{i}(s)]{\rm var}[X_{j}(s)]}}=\frac{\beta_{ij}}{\sqrt{\beta_{ii}\beta_{jj}}}\frac{\left[a_{i}(s)a_{j}(s)\right]^{1/4}}{\sqrt{\alpha_{ij}(s,s)}},$
for each $s\in\mathcal{L}$ and $i\neq j$.
Using the representation in terms of generalized Stieltjes functions of order $1/2$, as described in Section 3.2, the function (4.2) can be expressed as
$K_{ij}(s,t)=\begin{cases}\displaystyle\beta_{ij}\sqrt{\frac{b_{i}(s)b_{j}(s)}{\alpha_{ij}(s,s)}}F((0,\infty)),&\text{ if }s=t,\\ \\ \displaystyle\beta_{ij}\sqrt{\frac{b_{i}(s)b_{j}(t)}{\text{d}_{R}(s,t)}}\psi\left(\frac{\alpha_{ij}(s,t)}{\text{d}_{R}(s,t)}\right),&\text{ if }s\neq t,\end{cases}$
with $\psi$ defined according to Equation (3.2). Thus, analogous examples to those presented in Section 3.2 can be constructed for the vector-valued case.
In particular, when $a_{i}$ and $b_{i}$ are constant functions for all $i=1,\ldots,q$, Equation (4.2) defines a family of matrix-valued isotropic covariance functions. Although this is a special case, it provides a useful multivariate extension of a subclass of the isotropic covariance models introduced by Anderes et al., (2020). It also complements previous results on vector-valued random fields on linear networks by Filosi et al., (2025), where isotropic matrix-valued covariance functions on trees were derived.
@@ 5.1 Description of the method
We consider the weighted local likelihood estimation proposed by Anderes and Stein, (2011). We now describe the method and the necessary modifications required to adapt it to the network context.
Let $X(s_{1})\ldots,X(s_{N})$ be a realization of a GRF on $\mathcal{L}$. For any given site $s\in\mathcal{L}$, let $\mathcal{N}_{s,k}$ denote the set of $k$ observations nearest to $s$ with respect to the resistance metric. Recall that our goal is to estimate $a$ and $b$ at the selected site $s$. Given weights $\lambda_{1},\ldots,\lambda_{N}$, the weighted local likelihood function is defined as
@@
where $\text{Lik}(a_{0},b_{0}\mid\mathcal{N}_{s,k})$ denotes the log-likelihood of the data in $\mathcal{N}_{s,k}$, with $a_{0}$ and $b_{0}$ representing fixed values of the functions $a(\cdot)$ and $b(\cdot)$. As in Anderes and Stein, (2011), we define $\text{Lik}(a_{0},b_{0}\mid\mathcal{N}_{s,0})$ as 0.
@@
The principle behind this method is that we locally fit an isotropic model, not simply by taking a window of data around $s$ (a strategy also used in Anderes and Stein,, 2011, which has been shown to be less competitive), but by considering inflating neighborhoods and gradually assigning less weight to those that cover a very large region. For this reason, the weights include a bandwidth parameter to control this effect.
5.2 Simulation
@@
$\lambda_{k}^{\text{cs}}=\left(1-\frac{\text{d}_{R}(s,s_{k})}{\tau}\right)_{+},\qquad k=1,\ldots,N,$ (5.1)
@@
$\lambda_{k}^{\text{g}}=\exp\left(-\frac{\text{d}_{R}(s,s_{k})^{2}}{\eta^{2}}\right),\qquad k=1,\ldots,N,$ (5.2)
@@
We consider the Chicago street network available in the R package spatstat, and the covariance function in Equation (3.3), with the parameterization discussed in Section 3.3. Following Anderes and Stein, (2011), we assume for simplicity that the variance of the random field is known and equal to one, so that we focus on the estimation of the local practical range. The function $a$ used in our study is shown in the left panel of Figure 5.1. We also highlight six randomly selected locations on the network at which $a$ will be estimated.
We simulate 100 independent realizations of the GRF on this domain at 503 locations (one at the midpoint of each edge). For each realization, we estimate $a$ at the six selected sites. The weighted local likelihood method is implemented with the weighting schemes described above. We use bandwidth parameters $\eta=120$ and $\tau=200$, which produce weight functions with a comparable spatial extent but distinct structural features (see the right panel of Figure 5.1). For reference, the maximum resistance distance between two points in this network is approximately 675.
Figure 5.1: (Left) The function $a$ used in our simulations on the Chicago street network. Six randomly selected locations where $a$ will be estimated are highlighted. (Right) Decay of the weights as a function of the resistance metric for the compactly supported ($\tau=120$) and Gaussian ($\eta=120$) weighting schemes.
In Figure 5.2, we show the centered boxplots of the estimates across the 100 realizations, for each weighting scheme and each of the six selected sites. Overall, no significant bias is observed in any of the boxplots, demonstrating the effectiveness of the method. One point of interest is the higher variance in the estimates at site 2. This effect is consistent with the site’s proximity to the boundary of the network, which reduces the number of neighbors that effectively contributes to the local estimate. A similar observation was reported by Anderes and Stein, (2011).
Figure 5.2: Centered boxplots of the estimates of $a$, at the six selected sites, across 100 independent realizations of a GRF with covariance function (3.3). Results are shown for the compactly supported and Gaussian weighting schemes.
@@
We also compare the computational performance of the weighting schemes by measuring evaluation times as a function of the sample size of the GRF. Specifically, we simulate GRFs with $N=503,1006,1509$, and 2012, which correspond to taking 1, 2, 3, and 4 locations per edge, respectively, and implement the method using compactly supported weights with bandwidth parameters $\tau=100,150$, and 200. For each configuration, we report the time required for a single evaluation of the weighted local likelihood function.
The results are summarized in Table 5.1, allowing us to assess the computational gains relative to the Gaussian weighting scheme. For $\tau=100$ and 150, the reduction in computation time is particularly pronounced. For instance, when $\tau=150$, the computation time is reduced by approximately $50\%$ compared with Gaussian weighting scheme, whereas for $\tau=100$, evaluating the objective function takes at most $2\%$ of the time required with Gaussian weights, and for most sample sizes the percentage is even lower.
Table 5.1: Times (in seconds) for a single evaluation of the weighted local likelihood function under each weighting scheme, as a function of the sample size NN. For the compactly supported weights we consider $\tau=100,150$, and 200. The simulation was performed on an Apple M1 chip with 8GB of memory.
@@
Our findings both support the observations reported in Anderes and Stein, (2011) regarding reduced precision near the domain boundaries and provide additional insights into the statistical and computational performance of different weighting schemes. A promising strategy is to use a combination of weights, employing compactly supported weights in the interior and adapting the weights near the boundaries. We refer the reader to Anderes and Stein, (2011) for a more comprehensive treatment of weights and bandwidth selection.